# Cavity-Altered Superconductivity

### v3

@@ Methods
Heterostructures were fabricated by placing vdW materials, on the order of 10 nm to 120 nm thick, on large flat surfaces of -ET single crystals (ac crystallographic plane, Supplementary Information Section 2). Measurements were performed on five devices (Supplementary Information Section 3). Data in Fig. 1 are from device 1, in Fig. 2 from device 2, in Fig. 3(b-d) from device 1, and in Fig. 3a from device 5. Fig. 4 is a compilation of results from devices 1, 3, and 4.
The -ET crystal surfaces were either as-grown (devices 1 and 3), exposed by cleaving bulk crystals (devices 4 and 5), or exposed by exfoliating single crystals (device 2). For bulk samples, the -ET single crystal was fixed onto a silicon substrate using either epoxy or silver paint. For devices 4 and 5, the crystal was then cleaved to expose new surfaces. Large atomically flat surfaces were then selected by optical and atomic force microscopy. As-grown surfaces were regularly found to be atomically flat. Cleaved surfaces showed larger variations in surface roughness.
Both hBN and RuCl3 flakes were exfoliated using a standard scotch-tape technique onto a polydimethylsiloxane (PDMS) stamp and identified via optical and atomic force microscopy. The stacking process was performed using a commercial transfer stage with temperature control. The PDMS stamp was gently pressed onto the -ET surface and then slowly retracted. The hBN flakes were released at approximately 60°C, while the RuCl3 flakes were released at around 110°C. Isotopically pure hBN (10BN) was used to achieve a narrow linewidth for the TO phonon. The narrow linewidth makes HPhPs visible in s-SNOM.
hBN/BSCCO samples were prepared by exfoliating and assembling BSCCO and hBN crystals in a glovebox filled with inert gas ( ppm; ppm). BSCCO was first exfoliated on SiO2 substrates. hBN flakes were then exfoliated on a PDMS stamp and transferred onto BSCCO at room temperature.
@@
To minimize noise and instrument drift, was typically averaged over multiple measurements repeated at each point and taken with high resolution (typically 1-2 nm to define the surface accurately). During the measurements, the tip was electrically biased to compensate for the local contact potential difference and thus minimize electrostatic forces. The average curves were then smoothed with a uniform filter with width 30 to 50 pixels to reduce the appearance of noise. For the -ET heterostructures, a linear background was subtracted by fitting the data far from the sample surface where the Meissner force is negligible. For the BSCCO sample, a background measured on the silicon substrate was subtracted (Supplementary Information Section 5). For proper comparison in Fig. 1, we have compensated for the thickness of the hBN and RuCl3 microcrystals such that the measurements for hBN/-ET, bare -ET, and RuCl3/-ET all have the -ET surface at zero.
@@
Acknowledgments We acknowledge helpful discussions with O. Auslaender. This work was supported by the European Research Council (ERC-2024-SyG- 101167294 ; UnMySt), the Cluster of Excellence Advanced Imaging of Matter (AIM), Grupos Consolidados (IT1249-19) and SFB925, We acknowledge support from the Max Planck-New York City Center for Non-Equilibrium Quantum Phenomena. The Flatiron Institute is a division of the Simons Foundation. Support for hexagonal boron nitride crystal growth was provided by the Office of Naval Research, Award No. N00014-22-1-2582. K.K. acknowledges the support of the JSPJ KANENHI (21K18144), the JSPS Core-to-Core Program (JPJSCCA20240001) and the Alexander von Humboldt Foundation. The work at BNL was supported by the US Department of Energy, office of Basic Energy Sciences, contract no. DOE-SC0012704. D.M. and M.C. acknowledge support from the Gordon and Betty Moore Foundation’s EPiQS Initiative, Grant GBMF9069. D.M.K. and M.A.S. acknowledge funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) - 531215165 (Research Unit ‘OPTIMAL’). M.A.S. was funded by the European Union (ERC, CAVMAT, project no. 101124492). T.O. acknowledges support from the JSPS Overseas Research Fellowships. D.S. was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. RS-2024-00333664). J. Y. and Q.L were supported by the Office of Basic Energy Sciences, Materials Sciences and Engineering Division, U.S. Department of Energy (DOE) under Contract No. DE-SC0012704. We acknowledge project/application support by the Max Planck Computing and Data Facility.
@@
Author contributions I.K., T.A.W., J.X., and G.P. collected and analyzed MFM data. I.K. and T.A.W. wrote the manuscript. S.Z. and S.S.Z. collected SNOM measurements. S.Z and J.Z. analyzed SNOM data. S.Z. simulated the HPhP dispersion for hBN on -ET. D.S. and B.S.Y.K. fabricated the heterostructures with input from C.R.D. G.P. performed SEM imaging of the MFM tips. J.Y. and Q.L. performed SQUID magnetometry measurements. T.O. provided simulation of the electric field in a z-dependent superfluid density scenario. J.H.E. grew the hBN. M.J., R.P., and S.W. provided discussion and interpretation of the data, as well as formulation of the idea. K.M. and K.K. grew the -ET crystals. G.G grew the BSCCO crystal. D.M. and M.C. grew the RuCl3 crystal. M.B. and A.C. provided discussion and interpretation of the data. D.M.K. provided guidance on fitting and participated in the discussion on mechanisms. D.S. performed the ab initio calculations under the supervision of A.R. D.S., M.H.M., and A.R. analyzed the ab initio results and developed a microscopic model for the quenching of superconductivity. M.A.S. provided theoretical input on the interpretation of results and contributed to the manuscript. A.N.P. advised on measurements and data interpretation. D.N.B. advised and supervised the project. All authors discussed the results and contributed to the final paper.

### v4

@@ Methods
The heterostructures were fabricated by placing vdW materials of thicknesses on the order of 10 nm to 120 nm, on large flat surfaces of -ET single crystals (ac crystallographic plane, Supplementary Information Section 2). Measurements were performed on six devices (Supplementary Information Section 4). Data in Fig. 1 are from Device 1, in Fig. 4 from Device 2, in Fig. 2(b-d) from Device 1, and in Fig. 2a from Device 6. The data in Fig. 3 represent a compilation of results from Devices 1, 3, 4, and 5.
The -ET crystal surfaces were either as-grown (Devices 1, 4, 5), exposed by cleaving bulk crystals (Devices 3 and 6), or exposed by exfoliating single crystals (Device 2). For bulk samples, the -ET single crystal was fixed onto a silicon substrate using either epoxy or silver paint. For Devices 3 and 6, the crystal was then cleaved to expose new surfaces. Large atomically flat surfaces were then selected by optical and atomic force microscopy. As-grown surfaces were regularly found to be atomically flat. Cleaved surfaces showed larger variations in surface roughness. For each device, MFM measurements spanned time periods ranging from weeks to months, and in general included multiple thermal cycles. For all devices, no sign of degradation was found.
Both hBN and RuCl3 microcrystals were exfoliated using a standard scotch-tape technique onto a polydimethylsiloxane (PDMS) stamp and identified via optical and atomic force microscopy. The stacking process was performed using a commercial transfer stage with temperature control. The PDMS stamp was gently pressed onto the -ET surface and then slowly retracted. The hBN microcrystals were released at approximately 60°C, while the RuCl3 microcrystals in Devices 1 and 5 were released at around 110°C and 130°C, respectively. Isotopically pure hBN (10BN) was used to achieve a narrow linewidth for the TO phonon. The narrow linewidth makes HPhPs visible in s-SNOM.
hBN/BSCCO samples were prepared by exfoliating and assembling BSCCO and hBN crystals in a glovebox filled with inert gas ( ppm; ppm). BSCCO was first exfoliated on SiO2 substrates. hBN microcrystals were then exfoliated on a PDMS stamp and transferred onto BSCCO at room temperature.
@@
To minimize noise and instrument drift, was typically averaged over multiple measurements repeated at each point and taken with high resolution (typically 1-2 nm to define the surface accurately). During the measurements, the tip was electrically biased to compensate for the local contact potential difference and thus minimize electrostatic forces. The average curves were then smoothed with a uniform filter with width 30 to 50 pixels to reduce the appearance of noise. For the -ET heterostructures, a linear background was subtracted by fitting the data far from the sample surface where the is negligible. For the BSCCO sample, a background measured on the silicon substrate was subtracted (Supplementary Information Section 6). For proper comparison in Fig. 1, we have compensated for the thickness of the hBN and RuCl3 microcrystals such that the measurements for hBN/-ET, bare -ET, and RuCl3/-ET all have the -ET surface at zero.
@@
Acknowledgments We acknowledge helpful discussions with O. Auslaender. The authors acknowledge the use of facilities and instrumentation supported by NSF through the Columbia University, Columbia Nano Initiative, and the Materials Research Science and Engineering Center DMR-2011738. D.N.B is a Moore Investigator in Quantum Materials EPIQS GBMF9455. Research at Columbia on electrodynamics of van der Waals interfaces is supported by DOE-BES DE-SC0018426. Fabrication of vdW heterostructures at Columbia and the development of nano-optical methods is supported as part of Programmable Quantum Materials, an Energy Frontier Research Center funded by the U.S. Department of Energy (DOE), Office of Science, Basic Energy Sciences (BES), under award DE-SC0019443. Advances in MFM imaging at Columbia are supported by ARO award W911NF2510062. This work was supported by the European Research Council (ERC-2024-SyG- 101167294 ; UnMySt), the Cluster of Excellence Advanced Imaging of Matter (AIM), Grupos Consolidados (IT1249-19) and SFB925, We acknowledge support from the Max Planck-New York City Center for Non-Equilibrium Quantum Phenomena. The Flatiron Institute is a division of the Simons Foundation. Support for hexagonal boron nitride crystal growth was provided by the Office of Naval Research, Award No. N00014-22-1-2582. K.K. acknowledges the support of the JSPJ KANENHI (21K18144), the JSPS Core-to-Core Program (JPJSCCA20240001) and the Alexander von Humboldt Foundation. The work at BNL was supported by the US Department of Energy, office of Basic Energy Sciences, contract no. DOE-SC0012704. D.M. and M.C. acknowledge support from the Gordon and Betty Moore Foundation’s EPiQS Initiative, Grant GBMF9069. D.M.K. and M.A.S. acknowledge funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) - 531215165 (Research Unit ‘OPTIMAL’). M.A.S. was funded by the European Union (ERC, CAVMAT, project no. 101124492). T.O. acknowledges support from the JSPS Overseas Research Fellowships. D.S. was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. RS-2024-00333664). J. Y. and Q.L were supported by the Office of Basic Energy Sciences, Materials Sciences and Engineering Division, U.S. Department of Energy (DOE) under Contract No. DE-SC0012704. We acknowledge project/application support by the Max Planck Computing and Data Facility.
@@
Author contributions I.K., T.A.W., J.X., and G.P. collected and analyzed MFM data. I.K. and T.A.W. wrote the manuscript. S.Z. and S.S.Z. collected SNOM measurements. S.Z and J.Z. analyzed SNOM data. S.Z. simulated the HPhP dispersion for hBN on -ET. D.S. and B.S.Y.K. fabricated the heterostructures with input from C.R.D. G.P. performed SEM imaging of the MFM tips. J.Y. and Q.L. performed SQUID magnetometry measurements. T.O. provided simulations of the electric field in a z-dependent superfluid density scenario, and in an oscillating dipole above hBN scenario. J.H.E. grew the hBN. M.J., R.P., and S.W. provided discussion and interpretation of the data, as well as formulation of the idea. K.M. and K.K. grew the -ET crystals. G.G grew the BSCCO crystal. D.M. and M.C. grew the RuCl3 crystal. M.B. and A.C. provided discussion and interpretation of the data. D.M.K. provided guidance on fitting and participated in the discussion on mechanisms. M.H.M. computed the quantum electric field fluctuations from a thin film of hBN. D.S. performed the ab initio calculations under the supervision of A.R. D.S., M.H.M., and A.R. analyzed the ab initio results and developed a microscopic model for the quenching of superconductivity. E.V.B. developed the analytical model for BCS-HM coupling. M.A.S. provided theoretical input on the interpretation of results and contributed to the manuscript. A.N.P. advised on measurements and data interpretation. D.N.B. advised and supervised the project. All authors discussed the results and contributed to the final paper.